# biblio.bib

### v1

@@ 1 Introduction
admits solutions. The critical value is also deeply connected to the homogenization of Hamilton–Jacobi equations, which starting from the seminal paper [LionsPapanicolaouVaradhan87] has received a considerable interest both from theoretical and from an applied viewpoint, see for example [Capuzzo-DolcettaIshii01, MajdaSouganidis94, ContrerasIturriagaSiconolfi14]. It then comes as no surprise that many works, such as [GomesOberman04, AchdouCamilliCapuzzo-Dolcetta08, CamilliCapuzzoDolcettaGomes07], are devoted to the approximation of the critical value for problems posed on $\mathds{T}^{N}$.
The aim of this paper is to approximate the critical value of eikonal equations posed on networks. With network we mean a connected finite graph $\Gamma$ embedded in $\mathds{R}^{N}$ with vertices linked by regular simple curves $\gamma$, called arcs of the network. A Hamiltonian on $\Gamma$ is a collection of Hamiltonians $\{H_{\gamma}\}$ indexed by the arcs, depending on state and momentum variables, with the crucial feature that Hamiltonians associated to arcs possessing different support are totally unrelated. The eikonal problem on $\Gamma$, as defined in [SiconolfiSorrentino18], is to find a continuous function $u:\Gamma\to\mathds{R}$ such that $u\circ\gamma$ is a viscosity solution to
@@
for any arc $\gamma$, where $c\in\mathds{R}$ is a fixed parameter. As in the torus case, the critical value of the eikonal equation is the unique value of the parameter $c$ for which (1) admits solutions and is deeply related to the so-called effective Hamiltonian of a corresponding homogenization problem, see [PozzaSiconolfiSorrentino24].
In general, (1) cannot be solved explicitly. Moreover, to design a numerical scheme for such an equation is a very difficult task since it requires the approximation of a first order Hamilton–Jacobi problem in which both the solution $u$ and the critical value $c$ are unknown. Indeed, there are few papers devoted to the numerical approximation of solutions to Hamilton–Jacobi equations on networks and most of them are focused on time dependent problems as described in [ImbertMonneau17, Siconolfi22], which are, from the numerical point of view, simpler than the eikonal one. Let us mention the finite difference based methods outlined in [CamilliCarliniMarchi18, CostesequeLebacqueMonneau14] and the semi-Lagrangian schemes in [CarliniFestaForcadel20, CarliniSiconolfi23, CamilliFestaSchieborn13].
Recently, in [Pozza23], it has been shown that the large time behavior of solutions to time-dependent Hamilton–Jacobi equations can be exploited to obtain solutions to a corresponding eikonal problem. Accordingly, our idea is to use the large time behavior of such solutions to approximate the critical value. The main difficulty of this approach is that the analysis carried out in [Pozza23] relies on a condition, denoted here by (D), which essentially requires knowing beforehand the value of $c$. This assumption is clearly too restrictive for our case, therefore we overcome it in this paper.
We provide two numerical schemes for the approximation of the critical value. The first one, Algorithm 1, uses an approach already employed in [AchdouCamilliCapuzzo-Dolcetta08] for problems posed on $\mathds{T}^{N}$ and allows an a priori error estimate. For Algorithm 2 we provide a convergence theorem and our simulations show that it usually achieves the same results of Algorithm 1 with fewer steps.
The paper is organized as follows: Section 2 provides some basic facts about networks and Hamiltonians defined on them, and gives our main assumptions. In Section 3 we introduce eikonal and time dependent equations, together with some results relevant to our study. Section 4 is about the analysis of our approximation strategies for the critical value. Finally, in Section 5, we present the numerical simulations of our algorithms. Appendix A is devoted to the proof of a Theorem which extends the large time behavior analysis performed in [Pozza23] to our framework.
@@ Acknowledgments.
The first author was partially supported by Italian Ministry of Instruction, University and Research (MIUR) (PRIN Project2022238YY5, “Optimal control problems: analysis, approximation”) and INdAM-research group GNCS (CUP__\__ E53C23001670001, “Metodi numerici innovativi per equazioni di Hamilton–Jacobi”). The second author is a member of the INdAM research group GNAMPA.
@@ 2.1 Basic Definitions
where $\mathcal{E}$ is a finite collection of regular (i.e., $C^{1}$ with non-vanishing derivative) simple oriented curves, called arcs of the network, with Euclidean length $|\gamma|$ and parameterized by arc length in $[0,|\gamma|]$ (i.e., $|\dot{\gamma}|\equiv 1$ for any $\gamma\in\mathcal{E}$). Note that we are also assuming existence of one-sided derivatives at the endpoints 0 and $|\gamma|$. We stress out that a regular change of parameters does not affect our results.
Observe that on the support of any arc $\gamma$, we also consider the inverse parametrization defined as
@@
$\gamma((0,|\gamma|))\cap\gamma^{\prime}\left(\left[0,\left|\gamma^{\prime}\right|\right]\right)=\emptyset,\qquad\text{whenever }\gamma^{\prime}\neq\gamma,\widetilde{\gamma}.$ (2)
@@
$\gamma((0,|\gamma|))\cap\mathbf{V}=\emptyset,\qquad\text{for any }\gamma\in\mathcal{E}.$
@@
This condition is due to the fact that in the known literature about time dependent Hamilton–Jacobi equations on networks no loops are admitted, see, e.g., [ImbertMonneau17, Siconolfi22, LionsSouganidis17].
@@ 2.2 Hamiltonians on $\Gamma$
[…]
@@
    If condition (H4) is satisfied, the above assumption is equivalent to
    $\operatorname{int}\{\mu\in\mathds{R}:H_{\gamma}(s,\mu)\leq a\}=\{\mu\in\mathds{R}:H_{\gamma}(s,\mu)<a\},\qquad\text{for any }a\in\mathds{R},\,s\in[0,1],$
    where $\operatorname{int}$ denotes the interior of a set.
@@
Under conditions (H1), (H3) and (H4) it is natural to define, for any $\gamma\in\mathcal{E}$, the Lagrangian corresponding to $H_{\gamma}$ as
@@ 3 Hamilton–Jacobi Equations on Networks
3.1 Eikonal HJ Equations and Critical Values
Here we are interested in equations, thoroughly analyzed in [SiconolfiSorrentino18, Pozza24], of the form
[…]
where $a\in\mathds{R}$. This notation synthetically indicates the family of Hamilton–Jacobi equations
[…]
for $\gamma$ varying in $\mathcal{E}$.
[…]
Definition 3.1.
[…]
  1. i)
[…]
  1. ii)
[…]
  2. iii)
    for every vertex $x$ there is at least one arc $\gamma\in\Gamma_{x}$ such that
    $H_{\gamma}(0,\partial_{s}\varphi(0))\geq a$
    for any $\varphi$ that is a constrained $C^{1}$ subtangent to $v\circ\gamma$ at 0.
[…]
We proceed to list some results about this problem, taken from [Pozza24], useful to our analysis. We start defining
$a_{\gamma}\coloneqq\max_{s\in[0,|\gamma|]}\min_{\mu\in\mathds{R}}H_{\gamma}(s,\mu),\quad\text{for each }\gamma\in\mathcal{E},\qquad a_{0}\coloneqq\max_{\gamma\in\mathcal{E}}a_{\gamma}$
and the critical value, or Mañé critical value, as
[…]
Now let $\sigma_{c}$ be a real map on the tangent bundle $T\Gamma$ such that:
  • if $x\in\mathbf{V}$ then
This map is crucial for the definition of the following uniqueness set:
Definition 3.2.
We call Aubry set on $\Gamma$, the closed set $\mathcal{A}_{\Gamma}$ made up of
[…]
We point out that $\mathcal{A}_{\Gamma}$ is not empty.
Theorem 3.3.
[…]
We conclude this Section by defining the semidistance on $\Gamma$
$S_{c}(y,x)\coloneqq\min\left\{\int_{0}^{T}\sigma_{c}\left(\xi,\dot{\xi}\right)\mathrm{d}\tau:\text{$\xi:[0,T]\to\Gamma$ is a curve from $y$ to }x\right\},$
whose importance is highlighted by the next Theorem, proved in [Pozza24, Theorem 4.8 and Corollary 4.10].
Theorem 3.4.
Let $\Gamma^{\prime}$ be a closed subset of $\Gamma$, $w:\Gamma^{\prime}\to\mathds{R}$ be a continuous function and define
[…]
3.2 Time Dependent HJ Equations
[…]
Here we focus on the following time dependent problem on $\Gamma$:
@@ 3.2 Time Dependent HJ Equations
Following [ImbertMonneau17], we call flux limiter any function $x\mapsto c_{x}$ from $\mathbf{V}$ to $\mathds{R}$ satisfying
@@
$c_{x}\coloneqq\max_{\gamma\in\Gamma_{x}}a_{\gamma},\qquad\text{for any }x\in\mathbf{V}.$ (5)
The definition of (sub/super)solutions to ($\mathcal{H}$JE) given in [PozzaSiconolfi23, Siconolfi22] is as follows:
Definition 3.5.
We say that a continuous function $w:\Gamma\times\mathds{R}^{+}\to\mathds{R}$ is a viscosity subsolution to ($\mathcal{H}$JE) if
  1. i)
[…]
  2. ii)
    for any $T\in(0,\infty)$ and vertex $x$, if $\psi$ is a $C^{1}$ supertangent to $w(x,\cdot)$ at $T$ then $\partial_{t}\psi(T)\leq-c_{x}$.
A continuous function $v:\Gamma\times\mathds{R}^{+}\to\mathds{R}$ is called a viscosity supersolution to ($\mathcal{H}$JE) if
  1. iii)
[…]
  2. iv)
    for every vertex $x$ and $T\in(0,\infty)$, if $\psi$ is a $C^{1}$ subtangent to $v(x,\cdot)$ at $T$ such that $\partial_{t}\psi(T)<-c_{x}$, then there is a $\gamma\in\Gamma_{x}$ such that
    $\partial_{t}\varphi(0,T)+H_{\gamma}(0,\partial_{s}\varphi(0,T))\geq 0$
    for any $\varphi$ that is a constrained $C^{1}$ subtangent to $(s,t)\mapsto v(\gamma(s),t)$ at $(0,T)$. We stress out that this condition does not require the existence of constrained subtangents.
Furthermore, we say that $u:\Gamma\times\mathds{R}^{+}\to\mathds{R}$ is a viscosity solution to ($\mathcal{H}$JE) if it is both a viscosity subsolution and supersolution.
We have a result concerning the existence of solutions.
Theorem 3.6.
[Siconolfi22, Proposition 7.8] Given an initial datum $\phi\in C(\Gamma)$, ($\mathcal{H}$JE) admits a unique solution $v$ such that $v(0,x)=\phi(x)$ for every $x\in\Gamma$. This solution is in addition uniformly continuous.
In view of the previous Theorem we define, for every $t\in\mathds{R}^{+}$, the nonlinear operator $\mathcal{S}(t)$ on $C(\Gamma)$ such that, for each $\phi\in C(\Gamma)$, $\mathcal{S}(t)\phi$ is the unique solution at the time $t$ to ($\mathcal{H}$JE) with initial datum $\phi$. The family of operators $\{\mathcal{S}(t)\}_{t\in\mathds{R}^{+}}$ form a semigroup whose main properties are summarized below.
Proposition 3.7.
[PozzaSiconolfi23, Proposition 4.13]
  1. i)
  2. ii)
  3. iii)
We will provide a representation formula for solution to ($\mathcal{H}$JE) using a Lagrangian defined on the whole tangent bundle $T\Gamma$ of the network, namely the map $L:T\Gamma\to\mathds{R}$ such that
@@
Following [PozzaSiconolfi23], the operators $\mathcal{S}(t)$ can then be represented through the integral formula
$(\mathcal{S}(t)\phi)(x)=\min\left\{\phi(\xi(0))+\int_{0}^{t}L\left(\xi,\dot{\xi}\right)\mathrm{d}\tau:\text{$\xi$ is a curve with }\xi(t)=x\right\}.$
We stress out that there exists an optimal curve for $(\mathcal{S}(t)\phi)(x)$.
@@
[Siconolfi22, Theorem 7.4] If the initial datum $\phi$ is Lipschitz continuous, then $(x,t)\mapsto(\mathcal{S}(t)\phi)(x)$ is Lipschitz continuous in $\Gamma\times\mathds{R}^{+}$.
3.3 Large Time Behavior
We preliminarily review the condition
  1. (D)
    for any $\gamma\in\mathcal{E}$ with $a_{\gamma}=c=a_{0}$ the map $s\mapsto\min\limits_{p\in\mathds{R}}H_{\gamma}(s,p)$ is constant in $[0,|\gamma|]$.
@@ 3.3 Large Time Behavior
We proceed as in [Pozza23], extending the Aubry set defined in Definition 3.2.
@@
Remark 3.10.
If (D) holds or $c>a_{0}$, then $\widetilde{\mathcal{A}}_{\Gamma}$ corresponds to the extended Aubry set defined in [Pozza23] and, since the flux limiter satisfies (5), to the Aubry set defined in [SiconolfiSorrentino18]. Conversely, if $c=a_{0}$ and (D) does not hold, even if the flux limiter is minimal $\mathcal{A}_{\Gamma}$ and $\widetilde{\mathcal{A}}_{\Gamma}$ are generally different.
The large time behavior of solutions to ($\mathcal{H}$JE) in our setting is described by the next Theorem, whose proof is given in Appendix A.
@@
[…]
@@
on $\widetilde{\mathcal{A}}_{\Gamma}$, as well as the maximal subsolution in $\Gamma$ equal to $w$ on $\widetilde{\mathcal{A}}_{\Gamma}$.
Let us now list some consequences of Theorem 3.11 which will provide the theoretical basis for the algorithms in Section 4.
Proposition 3.12.
@@
$(\mathcal{S}(t)\phi)(x)+ct\leq u(x)+\max_{y\in\Gamma}(\phi(y)-u(y)),\qquad\text{for any }(x,t)\in\Gamma\times\mathds{R}^{+}.$
@@
$(\mathcal{S}(t)\phi)(x_{1})+ct\leq\phi(x_{1}),\qquad\text{for any }t\in\mathds{R}^{+},$ (8)
thus Theorem 3.8 yields, for any $(x,t)\in\Gamma\times\mathds{R}^{+}$,
@@
Exploiting once again Propositions A.1 and 3.7 we get
$(\mathcal{S}(t)\phi)(x)+ct\geq u(x)+\min_{y\in\Gamma}(\phi(y)-u(y)),\qquad\text{for any }(x,t)\in\Gamma\times\mathds{R}^{+},$
@@
[…]
Proposition 3.14.
@@
Inequalities (11) and (12) show that the sequences $\{\widetilde{c}_{k}\}$ and $\{\undertilde{c}_{k}\}$ are nonincreasing and nondecreasing, respectively. Finally, we exploit Theorem 3.11 to obtain
@@
=\;= $\lim_{k\to\infty}\max_{x\in\Gamma}\frac{(\mathcal{S}((k-1)T)\phi)(x)+c(k-1)T-(\mathcal{S}(kT)\phi)(x)-ckT}{T}+c=c,$
@@
Following Proposition 3.14, we get that $\undertilde{c}_{k}\leq c\leq\widetilde{c}_{k}$ for any $k\in\mathds{N}$, which implies the next result.
Lemma 3.15.
Under the same assumptions of Proposition 3.14, we define for every $k\in\mathds{N}$
@@
4 Algorithms for the Critical Values
In this Section we present two algorithms for the approximation of the critical value. They require a method to approximate solutions to ($\mathcal{H}$JE), since they rely on the large time behavior of said solutions. We will employ the semi-Lagrangian scheme presented in [CarliniSiconolfi23]. It takes as inputs a Lipschitz continuous initial datum $\phi$, a final time $T>0$ and a pair $\Delta\coloneqq(\Delta x,\Delta t)$ so that

### v2

# Numerical Approximation of the Critical Value of Eikonal Hamilton–Jacobi Equations on Networks

@@ 1 Introduction
admits solutions. The critical value is also deeply connected to the homogenization of Hamilton–Jacobi equations, which starting from the seminal paper [19] has received a considerable interest both from theoretical and from an applied viewpoint, see for example [9, 21, 13, 15, 22]. It then comes as no surprise that many works, such as [16, 1, 5, 4], are devoted to the approximation of the critical value for problems posed on $\mathds{T}^{N}$.
The aim of this paper is to approximate the critical value of eikonal equations posed on networks. With network we mean a connected finite graph $\Gamma$ embedded in $\mathds{R}^{N}$ with vertices linked by regular simple curves $\gamma$, called arcs of the network. A Hamiltonian on $\Gamma$ is a collection of Hamiltonians $\{H_{\gamma}\}$ indexed by the arcs, depending on state and momentum variables, with the crucial feature that Hamiltonians associated to arcs possessing different support are totally unrelated. Over the last years there has been an increasing interest in the study of this type of problem, leading to many valuable contributions. Among these we cite [3, 8, 17, 18, 20, 2].
The eikonal problem on $\Gamma$, as defined in [28], is to find a continuous function $u:\Gamma\to\mathds{R}$ such that $u\circ\gamma$ is a viscosity solution to
@@
for any arc $\gamma$, where $c\in\mathds{R}$ is a fixed parameter. As in the torus case, the critical value of the eikonal equation is the unique value of the parameter $c$ for which (1) admits solutions and is deeply related to the so-called effective Hamiltonian of a corresponding homogenization problem, see [26].
In general, ˜1 cannot be solved explicitly. Moreover, to design a numerical scheme for such an equation is a very difficult task since it requires the approximation of a first order Hamilton–Jacobi problem in which both the solution $u$ and the critical value $c$ are unknown. Indeed, there are few papers devoted to the numerical approximation of solutions to Hamilton–Jacobi equations on networks and most of them are focused on time-dependent problems as described in [17, 27], which are, from the numerical point of view, simpler than the eikonal one. Let us mention the finite difference based methods outlined in [6, 14] and the semi-Lagrangian schemes in [10, 12, 7].
Recently, in [24], it has been shown that the large time behavior of solutions to time-dependent Hamilton–Jacobi equations can be exploited to obtain solutions to a corresponding eikonal problem. Accordingly, our idea is to use the large time behavior of such solutions to approximate the critical value.
We provide two numerical schemes for the approximation of the critical value. The first one, Algorithm˜1, uses an approach already employed in [1] for problems posed on $\mathds{T}^{N}$ and allows an a priori error estimate. For Algorithm˜2 we provide a convergence theorem and our simulations show that it usually achieves the same results of Algorithm˜1 with fewer steps.
The paper is organized as follows: Section˜2 provides some basic facts about networks and Hamiltonians defined on them, and gives our main assumptions. In Section˜3 we introduce the time-dependent equation and its large time behavior, which will provide the theoretical basis for our algorithms. Section˜4 is about the analysis of our approximation strategies for the critical value. Finally, in Section˜5, we present the numerical simulations of our algorithms. Appendix˜A is devoted to the proof of a Theorem which extends the large time behavior analysis performed in [24] to our framework.
@@ Acknowledgments.
The first author was partially supported by Italian Ministry of Instruction, University and Research (MIUR) (PRIN Project2022238YY5, “Optimal control problems: analysis, approximation”) by INdAM–GNCS Project (CUP_\_ E53C24001950001, “Metodi avanzati per problemi di Mean Field Games ed applicazioni”), by Avvio alla Ricerca (CUP_\_ B83C24006550001, “Numerical schemes for Hamilton-Jacobi equations and Mean Field Games problems”) and by European Union - Next Generation EU, Missione 4, Componente 1, CUP_\_ B53C23002010006. The second author is a member of the INdAM research group GNAMPA.
@@ 2.1 Basic Definitions
where $\mathcal{E}$ is a finite collection of regular (i.e., $C^{1}$ with non-vanishing derivative) simple oriented curves, called arcs of the network, with Euclidean length $|\gamma|$ and parameterized by arc length in $0,|\gamma|]$ (i.e., $|\dot{\gamma}|\equiv 1$ for any $\gamma\in\mathcal{E}$). Note that we are also assuming existence of one-sided derivatives at the endpoints 0 and $|\gamma|$. We stress out that a regular change of parameters does not affect our results.
On the support of any arc $\gamma$, we also consider the inverse parametrization defined as
@@
[…]
@@
[…]
@@
This condition is due to the fact that in the known literature about time-dependent Hamilton–Jacobi equations on networks no loops are admitted, see, e.g., [17, 27, 20]. We believe that our approach, the same used in [27], can also include the presence of loops, but this should require nontrivial adjustments like the addition of a periodic condition on loops.
@@ 2.2 Hamiltonians on $\Gamma$
$H_{\gamma}:[0,|\gamma|]\times\mathds{R}$ $\>\longrightarrow\>$ $\mathds{R}$
$(s,\mu)$ $\>\longmapsto\>$ $H_{\gamma}(s,\mu),$
@@
Under conditions˜(H1), (H3) and (H4) it is natural to define, for any $\gamma\in\mathcal{E}$, the Lagrangian corresponding to $H_{\gamma}$ as
@@ 3 Hamilton–Jacobi Equations on Networks
In this Section we briefly recall some results about Hamilton–Jacobi equations posed on networks which will form the theoretical basis of our algorithms.
3.1 Time-Dependent HJ Equations
We focus on the time-dependent problem, thoroughly analyzed in [27, 17], of the form
@@ 3.1 Time-Dependent HJ Equations
[…]
$\left\{(x,t):t\in\mathds{R}^{+}\right\}\qquad\text{with $x\in\mathbf{V}$}.$
Following [17], we set
$a_{\gamma}\coloneqq\max_{s\in[0,|\gamma|]}\min_{\mu\in\mathds{R}}H_{\gamma}(s,\mu),\quad\text{for each }\gamma\in\mathcal{E},$
and call flux limiter any function $x\mapsto c_{x}$ from $\mathbf{V}$ to $\mathds{R}$ satisfying
@@
$c_{x}\coloneqq\max_{\gamma\in\Gamma_{x}}a_{\gamma},\qquad\text{for any }x\in\mathbf{V}.$
The flux limiter, roughly speaking, is a choice of a constant on each vertex which bound from above the time derivative of any subsolutions on it. We refer to [27] for the definition of solution to $\mathcal{H}$JE and the role of the flux limiter in it.
Next we define a global Lagrangian $L:T\Gamma\to\mathds{R}$ by gluing together the $L_{\gamma}$ and taking into account the flux limiter $c_{x}$:
@@
[25, Theorem 6.7] Given an initial datum $\phi\in C(\Gamma)$, the function
@@
is the unique solution to $\mathcal{H}$JE with $\mathcal{S}(0)\phi=\phi$ and flux limiter $c_{x}$.
We stress out that there exists an optimal curve for $(\mathcal{S}(t)\phi)(x)$. Furthermore, the family of operators $\{\mathcal{S}(t)\}_{t\in\mathds{R}^{+}}$ form a semigroup whose main properties are summarized below.
Proposition 3.2.
[24, Proposition 4.13]
  1. i)
  2. ii)
  3. iii)
Theorem 3.3.
[27, Theorem 7.4] If the initial datum $\phi$ is Lipschitz continuous, then $(x,t)\mapsto(\mathcal{S}(t)\phi)(x)$ is Lipschitz continuous in $\Gamma\times\mathds{R}^{+}$.
3.2 Approximation of the Critical Value
We consider the eikonal equation
$\mathcal{H}(x,\mathrm{D}u)=a,\qquad\text{on }\Gamma,$ ($\mathcal{H}$Jaa)
namely the family of Hamilton–Jacobi equations
$H_{\gamma}(s,\partial U)=a,\qquad\text{on }[0,|\gamma|],$
where $a\in\mathds{R}$, for $\gamma$ varying in $\mathcal{E}$ and coupled with a state-constraint-type boundary condition at the vertices. The critical value, or Mañé critical value, is characterized in [28, 23] as the unique value
$c\geq a_{0}\coloneqq\max_{\gamma\in\mathcal{E}}a_{\gamma}$ (5)
for which ($\mathcal{H}$Jcc) (namely the equation ($\mathcal{H}$Jaa) with $a=c$) admits solutions.
It has been shown in [24] that $\mathcal{S}(t)\phi+ct$ converges to a solution to ($\mathcal{H}$Jcc) as $t\to\infty$. We will use this fact to approximate the critical value.
The large time behavior of the solutions to $\mathcal{H}$JE is mainly due to the relation between the optimal curves of ˜4 and the so called Aubry set. More in detail, let $\sigma_{c}$ be a real map on the tangent bundle $T\Gamma$ such that:
  • if $x\in\mathbf{V}$ then
Definition 3.4.
We call Aubry set on $\Gamma$, the closed set $\mathcal{A}_{\Gamma}$ made up of
  • the $x\in\Gamma$ incident to a closed curve $\xi:[0,T]\to\Gamma$ with $\int_{0}^{T}\sigma_{c}\left(\xi,\dot{\xi}\right)d\tau=0$ and a.e. non-vanishing derivative;
Setting the semidistance
$S_{c}(y,x)\coloneqq\min\left\{\int_{0}^{T}\sigma_{c}\left(\xi,\dot{\xi}\right)\mathrm{d}\tau:\text{$\xi:[0,T]\to\Gamma$ is a curve from $y$ to $x$}\right\}$
we further have
Theorem 3.5.
[23, Theorem 4.8 and Corollary 4.10] Let $\Gamma^{\prime}$ be a closed subset of $\Gamma$, $w:\Gamma^{\prime}\to\mathds{R}$ be a continuous function and define
Then $u$ is the maximal subsolution to ($\mathcal{H}$Jcc) not exceeding $w$ on $\Gamma^{\prime}$ and a solution in $\Gamma\setminus(\Gamma^{\prime}\setminus\mathcal{A}_{\Gamma})$. If $w$ is a subsolution $u$ agrees with it on $\Gamma^{\prime}$.
As pointed out in [24], the convergence of $\mathcal{S}(t)\phi+ct$ is influenced by the flux limiter. We account for this by extending, in a certain sense, the Aubry set.
@@ 3.2 Approximation of the Critical Value
The large time behavior of solutions to $\mathcal{H}$JE in our setting is described by the next Theorem, whose proof is given in Appendix˜A.
@@
The limit function ˜6 is a solution to ($\mathcal{H}$Jcc) in $\Gamma\setminus\left(\widetilde{\mathcal{A}}_{\Gamma}\setminus\mathcal{A}_{\Gamma}\right)$ agreeing with
@@
on $\widetilde{\mathcal{A}}_{\Gamma}$, as well as the maximal subsolution in $\Gamma$ equal to $w$ on $\widetilde{\mathcal{A}}_{\Gamma}$, see Theorem˜3.5.
Let us now list some consequences of Theorem˜3.7 which will provide the theoretical basis for the algorithms in Section˜4.
Theorem 3.8.
@@
Let $u$ be a solution to ($\mathcal{H}$Jcc) in $\Gamma\setminus\widetilde{\mathcal{A}}_{\Gamma}$. We know from Propositions˜A.1 and 3.2 that
@@
[…]
thus Theorem˜3.3 yields, for any $(x,t)\in\Gamma\times\mathds{R}^{+}$,
@@
Exploiting once again Propositions˜A.1 and 3.2 we get
[…]
@@
Next we provide a different approach to the approximation of the critical value. The main difference between Theorem˜3.8 and the Theorem below is that the former provides an estimate for the rate of convergence of $(\phi-\mathcal{S}(t)\phi)/t$ to $c$ as $t$ positively diverges, while the latter exploits the convergence of $\mathcal{S}(t)\phi+ct$ to a solution to ($\mathcal{H}$Jcc) to define two sequences converging monotonically to $c$.
Theorem 3.10.
@@
Inequalities (11) and (12) show that the sequences $\{\widetilde{c}_{k}\}$ and $\{\undertilde{c}_{k}\}$ are nonincreasing and nondecreasing, respectively. Finally, we exploit Theorem˜3.7 to obtain
@@
$=\;$ $\lim_{k\to\infty}\max_{x\in\Gamma}\frac{(\mathcal{S}((k-1)T)\phi)(x)+c(k-1)T-(\mathcal{S}(kT)\phi)(x)-ckT}{T}+c=c,$
@@
Following Theorem˜3.10, we get that $\undertilde{c}_{k}\leq c\leq\widetilde{c}_{k}$ for any $k\in\mathds{N}$, which implies the next result.
Corollary 3.11.
Under the same assumptions of Theorem˜3.10, we define for every $k\in\mathds{N}$
@@
4 Algorithms for the Critical Value
In this Section we present two algorithms for the approximation of the critical value.
4.1 Numerical Approximation of Time-Dependent HJ Equations
Our algorithms require a method to approximate solutions to ($\mathcal{H}$JE), since they rely on the large time behavior of said solutions.
We employ the semi-Lagrangian scheme presented in [12] to approximate the solutions to ($\mathcal{H}$JE). Rather than directly discretizing equation ($\mathcal{H}$JE), the scheme addresses a modified problem that involves modified Hamiltonians $\widetilde{H}_{\gamma}$ obtained from the $H_{\gamma}$ through the procedure described in [12, Appendix A]. This construction depends on the preliminary choice of a compact interval $I$ of the momentum variable $\mu$ within which the modified Hamiltonians coincide with the original ones:
$\widetilde{H}_{\gamma}(s,\mu)=H_{\gamma}(s,\mu),\qquad\text{for every }\gamma\in\mathcal{E},\,s\in[0,|\gamma|],\,\mu\in I.$
The key advantage of this approach is the existence of a positive constant $\beta_{0}$, depending on $I$, such that the modified Lagrangians
$\widetilde{L}_{\gamma}(s,\lambda)\coloneqq\max_{\mu\in\mathds{R}}\left(\lambda\mu-\widetilde{H}_{\gamma}(s,\mu)\right),\qquad\text{for }\gamma\in\mathcal{E},\,s\in[0,|\gamma|],\,\lambda\in\mathds{R},$
are equal to $\infty$ when $\lambda$ is outside $\beta_{0},\beta_{0}]$, for all $s\in[0,|\gamma|]$, $\gamma\in\mathcal{E}$, and are Lipschitz continuous in $\beta_{0},\beta_{0}]$. Setting $\widetilde{\mathcal{H}}\coloneqq\{\widetilde{H}_{\gamma}\}_{\gamma\in\mathcal{E}}$, the modified problem is
[…]
The scheme is based on a space-time discretization defined by a final time $T>0$ and a pair $\Delta\coloneqq(\Delta x,\Delta t)$ such that